# Preconditioned log-determinant approximation:

### v1

@@ 1 Introduction
This quantity arises in many applications, for instance, in machine learning and in statistics, when dealing with the discretization of covariance kernels, see e.g. [RasmussenWilliams06, Seeger04], and it plays a central role in the log-marginal likelihood for Gaussian processes [CoverThomas06, GardnerPleissWeinbergerBindelWilson18], which appears in hyperparameter optimization [WengerPleissHennigCunninghamGardner22] and in the normalization of the determinantal point processes for supervised learning [KuleszaTaskar12].
@@
which follows from Binet theorem. With a suitable choice of $P$, the term ${\rm trace}\log(P)$ can be computed cheaply and exactly, and the second term can be estimated via (SLQ). This idea has been explored, for instance, in [FengKulickTang24, WengerPleissHennigCunninghamGardner22]. A good preconditioner can be constructed as $P=S+I$, where $S$ is a low-rank approximation of $A$. In this sense, (low-rank) corresponds to the approximation of (2) in which the last term is completely ignored. A black-box preconditioner can be constructed by taking, as $S$, the Nyström approximation [HalkoMartinssonTropp11, LiLindermanSzlamStantonKlugerTygert17]:
@@ 1.1 Contributions
We show – in practice and in theory – that this strategy is effective when the matrix $A$ has at least moderate spectral decay. The advantage of our one-sample strategy over all possible instances of (2) is twofold: having a good preconditioner reduces the variance of the Girard-Hutchinson trace estimator, and the condition number of the matrix $\widehat{M}_{\ell}$ is small so Lanczos quadrature is expected to be precise with only a small number of steps.
@@ 1.2 Relation with existing literature
The idea of using a single sample for trace estimation problems involving the Girard-Hutchinson estimator has appeared before and has been proven effective for a number of problems. More specifically, it has been used for trace estimation of matrices which are available via matrix-vector multiplications (matvecs), XTrace [EpperlyTroppWebber24], in the context of randomized probing methods [FrommerRinelliSchweitzer25], and for spectral gap estimation problems [BenziRinelliSimunec24]. The idea of using only one sample for (preconditioned) log-determinant computations, to the best of our knowledge, is new. Let us briefly review existing strategies to combine SLQ with preconditioning for approximating log-determinants. In [WengerPleissHennigCunninghamGardner22], the budget of matvecs in the approximation (2) is split evenly between the computation of the preconditioner and SLQ in the numerical experiments, whereas the rank of the low-rank approximation is taken to be equal to the number of samples for the Girard-Hutchinson trace estimator in the theoretical analysis. Another line of work addresses the approximation of traces of functions of numerically low-rank matrices by relying on the identity ${\rm trace}f(A)={\rm trace}f(P)+{\rm trace}(f(A)-f(P))$, for some low-rank approximation $P$ of $A$; the paper [PerssonKressner23] proposes funNys++, for the case of matrix monotone functions; a “Krylov-aware” version of funNys++ has been proposed in [ChenHallman23]. In the case of $f(x)=\log(1+x)$, these methods still need to apply Lanczos quadrature to the original matrix $A$, while the advantage of (2) is that one applies Lanczos quadrature to the preconditioned matrix and, in general, fewer matvecs are needed.
@@ 2 Analysis of the proposed strategy in the idealized case
Moreover, we can consider a “mixed” strategy that uses the best rank-$\lfloor\alpha\ell\rfloor$ approximation of the matrix for some $\alpha\in[0,1]$ and $\left\lfloor\frac{\ell+m-\lfloor\alpha\ell\rfloor}{m}\right\rfloor$ samples for the Girard-Hutchinson estimator (for simplicity, from now on we assume that both $\lfloor\alpha\ell\rfloor$ and $\left\lfloor\frac{\ell+m-\lfloor\alpha\ell\rfloor}{m}\right\rfloor$ are integers), which is
@@
where we use $w_{I}$ as a shorthand for the set of random vectors $w_{1},w_{2},\ldots,w_{\frac{(1-\alpha)\ell+m}{m}}$. In expectation, we have
@@
Figure 1 shows a comparison of the average errors (8), (9), and (10) for six matrices, described in Section 5.1, for a budget of matvecs ranging from $110$ to $1010$ and $m=10$. Specifically, the plots show the square root of the error (8) in thick blue lines, the error (9) in thick orange lines, and the square root of the error (10) computed for values of $\alpha$ ranging from $0.1$ to $0.9$ in thin lines ranging from purple (0.1) to blue (0.9). All expected errors have been normalized by the target quantity ${\rm trace}\log(A+I)$.
Figure 1: Square root of the variance (8) for the (idealized one-sample) strategy divided by ${\rm trace}\log(A+I)$, error (9) for the (low-rank) strategy divided by ${\rm trace}\log(A+I)$ and square root of the variance (10) divided by ${\rm trace}\log(A+I)$ for (idealized $\alpha$-rank) strategies for $\alpha=0.1,\ldots,0.9$, represented by thin lines, shading from purple ($\alpha=0.1$) to blue ($\alpha=0.9$) considering $m=10$ and the matrices $A\in\mathbb{R}^{4000\times 4000}$ described in Section 5.1.
@@ 3.1 Analysis of the one-sample strategy
where the inequality follows by the Cauchy-Schwartz inequality, and the second equality is the variance of the Girard-Hutchinson estimator. We use the expressions $\mathbb{E}_{\Omega}$ and $\mathbb{E}_{w}$ to denote the fact that the expectation is taken with respect to the random variable $\Omega$ or $w$, respectively. To derive a bound for the right-hand side of (3.1), we need to recall some general results on the Nyström approximation. Since $A\succeq\widehat{A}_{\ell}\succeq 0$ (see, e.g., [GittensMahoney16, Lemma 1]), we have that $\widehat{P}_{\ell}=\widehat{A}_{\ell}+I\succeq I$. Moreover, $x\mapsto\log(1+x)$ is an operator monotone function, that is, a function that preserves the Löwner order of SPSD matrices, see [Chansangiam13, Example 3.6]; we thus have
@@
where the first inequality follows from the fact that we have $\|H\|_{F}\leq\|H\|^{\frac{1}{2}}_{*}\|H\|^{\frac{1}{2}}_{2}$ for any matrix $H$ and the second inequality follows from Cauchy-Schwartz. Let us bound the two terms appearing in the right-hand-side of (19) separately. We have
@@ 3.2 Comparison with low-rank and $\alpha$-rank
and using Lemma 3.1 with $\alpha\ell$ instead of $\ell$ we get, for any $k\geq 0$, $p\geq 2$ with $k+p=\alpha\ell$,
@@
The bound (25) is difficult to compare directly with the bounds (22) and (23), one reason being that they are true for any combination of $k$ and $p$ with fixed sum (and the optimal values also depend on the chosen method). While (8) predicts that the squared error of the ideal one-sample strategy is proportional to $\|\log(\bar{\Lambda}_{k}+I)\|_{F}^{2}$, introducing the Nyström approximation makes the bound for the one-sample strategy (22) proportional to $\|\log(\bar{\Lambda}_{k}+I)\|_{*}$. This makes the benefit of the one-sample strategy over the low-rank strategy less clear from a theoretical point of view. For these reasons, we postpone a more thorough comparison to the numerical experiments in Section 5, where we compare all three strategies on different examples, and we also compare the theoretical bounds for the one-sample strategy (22) and the low-rank strategy (25). We will show, empirically, that the upper bound (22) is often much smaller than (25).
@@ 3.3 Inexact quadratic forms with Lanczos method
If we were to use (27) to analyze our one-sample strategy, we would run into the issue that, when $w\sim\mathcal{N}_{0,I}$, we typically have $\|w\|\approx\sqrt{n}$. Therefore, (27) only predicts a small error for moderately large values of $m$, even in the presence of a well conditioned matrix $B$. However, in practice, we observed that a small value of $m$ is sufficient to get an approximation error which is small with respect to the overall error of the one-sample strategy; see Section 5 below.
@@ 4 Error control and a “detective” version
We propose to overcome the issue by checking, numerically, whether a Nyström approximation of high rank is really advantageous over one of smaller rank. More specifically, assume that we have a budget of $\ell+m$ matvecs. For a fixed parameter $\beta\in(0,1)$, assuming for simplicity that $\beta\ell$ and $\beta^{2}\ell$ are integer, we compare the rank-$\beta\ell$ Nyström approximation of $A$ with the one of rank-$\beta^{2}\ell$. If we are satisfied with the improvement of the quality of the low-rank approximation from rank $\beta^{2}\ell$ to rank $\beta\ell$, we apply the one-sample strategy: we enlarge the Nyström approximation using a total of $\ell$ matvecs and then use SLQ with one random Gaussian vector using $m$ Lanczos steps. Otherwise, we apply the ($\alpha$-rank) strategy with $\alpha=\beta$.
@@ 4.1 Evaluating the error of the Nyström approximation
(28) $\mathrm{err}^{2}_{F}\left(A,\beta\ell\right):=\frac{1}{\beta\ell}\sum_{i=1}^{\beta\ell}\left\lVert(A\Omega^{(i)})\omega_{i}\right\rVert^{2}\approx\left\lVert A-\widehat{A}_{\beta\ell}\right\rVert^{2}_{F},$
where $\Omega^{(i)}\in\mathbb{R}^{n\times(\beta\ell-1)}$ is the matrix $\Omega$ with the $i$-th column removed and $\omega_{i}$ is the $i$-th column of $\Omega$. Roughly speaking, this amounts to estimating the Frobenius norm of the $(\beta\ell-1)$-Nyström approximation by multiplication with one random vector, and averaging the estimate over all $\beta\ell$ choices of the vector which is “left out”. Once the factors of the rank-$\beta\ell$ Nyström approximation has been computed, the expression (28) can be evaluated by an additional $\mathcal{O}(\beta^{3}\ell^{3})$ operations; see [EpperlyTropp24, Algorithm 4.1].
@@ 4.2 Choice of $\beta$
$\mathbb{E}\left[\mathrm{err}_{\mathrm{\alpha R_{Nys}}}^{2}\left(A,\ell,\frac{3}{4},w\right)\right]=\frac{8m}{\ell+4m}\left\lVert\log\left(\widehat{M}_{\frac{3}{4}\ell}\right)\right\rVert_{F}^{2}\leq\frac{8m}{\ell+m}\|\log(A+I)\|_{F}^{2};$
@@ 5.1 Example matrices
The last synthetic matrix is built in such a way that we have control over all the gaps and all the clusters of the eigenvalues of the matrix $H$. It is inspired from [SorensenEmbree16], generalized with several gaps as
@@ 5.4 Choice of the parameter $\beta$
We show the behavior of the log-det-ective method (Algorithm 1) for different values of the initial budget allocated for the low-rank approximation of the matrix, represented by the value of $\beta$, while the number of steps used to approximate any quadratic form in Lanczos method is fixed to $m=10$. In the experiments we take $\beta\in\left\{\frac{1}{8},\frac{1}{4},\frac{1}{2},\frac{3}{4},\frac{7}{8}\right\}$ and values of $\ell$ ranging from $100$ to $1000$. In each plot, the $x$-axis is the total number of matvecs $\ell+m$ and the $y$-axis is the final relative error in the approximation of the log-determinant. For each value of $\ell$ that we consider, we run each algorithm $100$ times; the plots report the mean relative error, with respect to the target quantity ${\rm trace}\log(A+I)$, and the standard deviation obtained discarding the best and the worst $10\%$ errors over the runs, represented by vertical bars. The plots in Figure 6 show that if we choose a value of $\beta$ that is too small, the method can fail to detect when it is convenient to switch the strategy, and this is due to the fact that the rank-$\beta\ell$-approximation is very different from the rank-$\ell$ approximation. Specifically, in (Example Alg), and (Example Matérn) with $\nu=\frac{1}{2}$, Algorithm 1 with the values $\beta=\frac{1}{8}$, $\beta=\frac{1}{4}$ in (Example RBF) wrongly applies the one-sample strategy. Besides, in (Example RBF), values of $\beta=\frac{1}{8}$, $\beta=\frac{1}{4}$ and $\beta=\frac{1}{2}$ wrongly predicts the mixed strategy. From now on we fix $\beta=\frac{3}{4}$.

### v2

@@ 1 Introduction
This quantity arises in many applications: for instance, in machine learning and in statistics, when dealing with the discretization of covariance kernels, see e.g. [RasmussenWilliams06, Seeger04], and it plays a central role in the log-marginal likelihood for Gaussian processes [CoverThomas06, GardnerPleissWeinbergerBindelWilson18], which appears in hyperparameter optimization [WengerPleissHennigCunninghamGardner22] and in the normalization of the determinantal point processes for supervised learning [KuleszaTaskar12].
@@
which follows from the multiplicative property of the determinant. With a suitable choice of $P$, the term ${\rm trace}\log(P)$ can be computed cheaply and exactly, and the second term can be estimated via (SLQ). This idea has been explored, for instance, in [FengKulickTang24, WengerPleissHennigCunninghamGardner22]. A good preconditioner can be constructed as $P=S+I$, where $S$ is a low-rank approximation of $A$. In this sense, (low-rank) corresponds to the approximation of (2) in which the last term is completely ignored. A black-box preconditioner can be constructed by taking, as $S$, the Nyström approximation [HalkoMartinssonTropp11, LiLindermanSzlamStantonKlugerTygert17]:
@@ 1.1 Contributions
We show – in practice and in theory – that this strategy is effective when the matrix $A$ has at least moderate spectral decay. The advantage of our one-sample strategy over all possible instances of (2) is twofold: having a good preconditioner reduces the variance of the Girard-Hutchinson trace estimator, and the condition number of the matrix $\widehat{M}_{\ell}$ is small so Lanczos quadrature is expected to be precise with only a small number of steps, see Figure 5.
@@ 1.2 Relation with existing literature
Since the vanilla Girard-Hutchinson trace estimator exhibits slow convergence, a variety of variance reduction techniques that apply to any symmetric (or SPSD) matrix have been developed. Most notably, Hutch++ [MeyerMuscoWoodruff21] combines the Girard-Hutchinson estimator with randomized low-rank approximation, and an adaptive version (A-Hutch++) was developed in [PerssonCortinovisKressner22]. In the context of numerically low-rank matrices, these techniques have been extended and proved to be effective for matrix functions, relying on the identity ${\rm trace}f(A)={\rm trace}f(P)+{\rm trace}\left(f(A)-f(P)\right)$, for some low-rank approximation $P$ of $A$. For instance, the paper [PerssonKressner23] proposes funNys++ for the case of matrix monotone functions. For the specific case of the logarithm $f(x)=\log(1+x)$, these methods still need to apply the Lanczos quadrature to the original matrix $A$. This means that they generally compare unfavorably with methods that leverage (2) to apply Lanczos quadrature to the preconditioned matrix instead, as we illustrate in Figure 7.
The idea of “preconditioning” the problem of computing traces of matrix functions has been used, implicitly, in the Krylov-aware algorithm [ChenHallman23]; see also [ChenHuberLinZaid26]. Other methods, tailored to log-determinant computations, leverage the explicit preconditioner (2): for instance, this was used in [SaibabaAlexanderianIpsen17] and [PerssonKressner23] to justify ignoring the $\mathrm{trace}\log\left(P^{-1/2}(A+I)P^{-1/2}\right)$ and concentrating on finding a good matrix $P$. Other works combine the benefit of both variance reduction and preconditioning in this context: in [GardnerPleissWeinbergerBindelWilson18] the incomplete Cholesky preconditioner is used, in [WengerPleissHennigCunninghamGardner22] the authors analyzed the effectiveness of various preconditioners, and they proposed, in the numerical experiments, to split the budget of matvecs evenly between the computation of the preconditioner and SLQ; despite being very general, their analysis involves error bounds which depend on the size of the matrix $A$. In [FengKulickTang24], the authors choose $P$ as the Nyström preconditioner of $A$, but no choice of parameters for the budget allocation and no convergence analysis is proposed. In our case, the analysis we propose for the Nyström preconditioner is independent from the size of $A$.
Lastly, the idea of using a single sample for trace estimation problems involving the Girard-Hutchinson estimator has appeared before and has been proven to be effective for a number of problems. More specifically, it has been used for trace estimation of matrices which are available via matrix-vector multiplications, XTrace [EpperlyTroppWebber24], in the context of randomized probing methods [FrommerRinelliSchweitzer25], and for spectral gap estimation problems [BenziRinelliSimunec24]. The idea of using only one sample for (preconditioned) log-determinant computations, to the best of our knowledge, is new.
@@ 2 Analysis of the proposed strategy in the idealized case
Moreover, we can consider a “mixed” strategy that uses the best rank-$\lfloor\alpha\ell\rfloor$ approximation of the matrix for some $\alpha\in[0,1]$ and $\left\lfloor\frac{\ell+m-\lfloor\alpha\ell\rfloor}{m}\right\rfloor$ samples for the Girard-Hutchinson estimator (for simplicity, from now on we assume that both $\alpha\ell$ and $\frac{\ell+m-{\alpha\ell}}{m}$ are integers), which is
@@
where we use $w_{I}$ as a shorthand for the set of random vectors $w_{1},\ldots,w_{\frac{(1-\alpha)\ell+m}{m}}$. In expectation, we have
@@
Figure 1 shows a comparison of the average errors (8), (9), and (10) for six matrices, described in Section 5.1, for a budget of matvecs ranging from $110$ to $1010$ and $m=10$. Specifically, the plots show the square root of the error (8) in thick blue lines, the error (9) in thick orange lines, and the square root of the error (10) computed for values of $\alpha$ ranging from 0 (which gives the approximation (SLQ)) to $0.9$ in thin lines ranging from purple (0) to blue (0.9). All expected errors have been normalized by the target quantity ${\rm trace}\log(A+I)$.
Figure 1: Square root of the variance (8) for the (idealized one-sample) strategy divided by ${\rm trace}\log(A+I)$, error (9) for the (low-rank) strategy divided by ${\rm trace}\log(A+I)$ and square root of the variance (10) divided by ${\rm trace}\log(A+I)$ for (idealized $\alpha$-rank) strategies for $\alpha=0,\ldots,0.9$, represented by thin lines, shading from purple ($\alpha=0$) to blue ($\alpha=0.9$) considering $m=10$ and the matrices $A\in\mathbb{R}^{4000\times 4000}$ described in Section 5.1.
@@ 3.1 Analysis of the one-sample strategy
where the inequality follows by the Cauchy-Schwarz inequality, and the second equality is the variance of the Girard-Hutchinson estimator. We use the expressions $\mathbb{E}_{\Omega}$ and $\mathbb{E}_{w}$ to denote the fact that the expectation is taken with respect to the random variable $\Omega$ or $w$, respectively. To derive a bound for the right-hand side of (3.1), we need to recall some general results on the Nyström approximation. Since $A\succeq\widehat{A}_{\ell}\succeq 0$ (see, e.g., [GittensMahoney16, Lemma 1]), we have that $\widehat{P}_{\ell}=\widehat{A}_{\ell}+I\succeq I$. Moreover, $x\mapsto\log(1+x)$ is an operator monotone function, that is, a function that preserves the Löwner order of SPSD matrices, see [Chansangiam13, Example 3.6]; we thus have
@@
where the first inequality follows from the fact that we have $\|H\|_{F}\leq\|H\|^{\frac{1}{2}}_{*}\|H\|^{\frac{1}{2}}_{2}$ for any matrix $H$ and the second inequality follows from Cauchy-Schwarz. Let us bound the two terms appearing in the right-hand-side of (19) separately. We have
@@ 3.2 Comparison with low-rank and $\alpha$-rank
and using Lemma 3.1 with $\alpha\ell$ instead of $\ell$ we get, for any $k\geq 0$, $p\geq 2$ with $k+p=\alpha\ell$, and $\alpha\ell\geq 4$,
@@
The bound (25) is difficult to compare directly with the bounds (22) and (23), one reason being that they are true for any combination of $k$ and $p$ with fixed sum (and the optimal values also depend on the chosen method). While (8) predicts that the squared error of the ideal one-sample strategy is proportional to $\|\log(\bar{\Lambda}_{k}+I)\|_{F}^{2}$, introducing the Nyström approximation makes the quantity $\|\log(\bar{\Lambda}_{k}+I)\|_{*}$ appear in the bound for the one-sample strategy (22). This makes the benefit of the one-sample strategy over the low-rank strategy less clear from a theoretical point of view. For these reasons, we postpone a more thorough comparison to the numerical experiments in Section 5, where we compare all three strategies on different examples, and we also compare the theoretical bounds for the one-sample strategy (22) and the low-rank strategy (25). We will show, empirically, that the upper bound (22) is often much smaller than (25).
@@ 3.3 Inexact quadratic forms with Lanczos method
If we were to use (27) to analyze our one-sample strategy, we would run into the issue that, when $w\sim\mathcal{N}_{0,I}$, we typically have $\|w\|\approx\sqrt{n}$. Therefore, (27) only predicts a small error for values of $m$ which are proportional to $\log(n)$, even in the presence of a well conditioned matrix $B$. However, in practice, we observed that a small constant value of $m$, e.g. $m=10$, is sufficient to get an approximation error which is small with respect to the overall error of the one-sample strategy; see Section 5 below.
@@ 4 Error control and a “detective” version
We propose to overcome the issue by checking, numerically, whether a Nyström approximation of high rank is really advantageous over one of smaller rank. More specifically, assume that we have a budget of $\ell+m$ matvecs. For a fixed parameter $\beta\in(0,1)$, assuming for simplicity that $\beta\ell$ and $\beta^{2}\ell$ are integers, we compare the rank-$\beta\ell$ Nyström approximation of $A$ with the one of rank-$\beta^{2}\ell$. If we are satisfied with the improvement of the quality of the low-rank approximation from rank $\beta^{2}\ell$ to rank $\beta\ell$, we apply the one-sample strategy: we enlarge the Nyström approximation using a total of $\ell$ matvecs and then use SLQ with one random Gaussian vector using $m$ Lanczos steps. Otherwise, we apply the ($\alpha$-rank) strategy with $\alpha=\beta$.
@@ 4.1 Evaluating the error of the Nyström approximation
(28) $\mathrm{err}^{2}_{F}\left(A,\beta\ell\right):=\frac{1}{\beta\ell}\sum_{i=1}^{\beta\ell}\left\lVert(A-\widehat{A}_{\Omega^{(i)}})\omega_{i}\right\rVert^{2}\approx\left\lVert A-\widehat{A}_{\beta\ell}\right\rVert^{2}_{F},$
where $\Omega^{(i)}\in\mathbb{R}^{n\times(\beta\ell-1)}$ is the matrix $\Omega$ with the $i$-th column removed, $\widehat{A}_{\Omega^{(i)}}$ is the Nyström approximation that results from leaving the $i$-th column out of $\Omega$, and $\omega_{i}$ is the $i$-th column of $\Omega$. Roughly speaking, this amounts to estimating the Frobenius norm of the $(\beta\ell-1)$-Nyström approximation by multiplication with one random vector, and averaging the estimate over all $\beta\ell$ choices of the vector which is “left out”. Once the factors of the rank-$\beta\ell$ Nyström approximation have been computed, the expression (28) can be evaluated by an additional $\mathcal{O}(\beta^{3}\ell^{3})$ operations; see [EpperlyTropp24, Algorithm 4.1].
@@ 4.2 Choice of $\beta$
$\mathbb{E}\left[\mathrm{err}_{\mathrm{\alpha R_{Nys}}}^{2}\left(A,\ell,\frac{3}{4},w\right)\right]=\frac{8m}{\ell+4m}\mathbb{E}\left[\left\lVert\log\left(\widehat{M}_{\frac{3}{4}\ell}\right)\right\rVert_{F}^{2}\right]\leq\frac{8m}{\ell+m}\|\log(A+I)\|_{F}^{2};$
@@ 4.3 Condition for switching the strategy
Remark 4.1.
Checking the condition (31) makes it so that Algorithm 1 is more likely to suggest the strategy (one-sample) over the strategy ($\alpha$-rank), with respect to the (idealized) case in which we check (29), because of the inequality $\log(1+x)\leq x$.
@@ 5.1 Example matrices
The last synthetic matrix is built in such a way that we have control over all the gaps and all the clusters of the eigenvalues of the matrix $H$. It is inspired by [SorensenEmbree16], generalized with several gaps as
@@ 5.4 Choice of the parameter $\beta$
We show the behavior of the log-det-ective method (Algorithm 1) for different values of the initial budget allocated for the low-rank approximation of the matrix, represented by the value of $\beta$, while the number of steps used to approximate any quadratic form in Lanczos method is fixed to $m=10$. In the experiments we take $\beta\in\left\{\frac{1}{8},\frac{1}{4},\frac{1}{2},\frac{3}{4},\frac{7}{8}\right\}$ and values of $\ell$ ranging from $100$ to $1000$. In each plot, the $x$-axis is the total number of matvecs $\ell+m$ and the $y$-axis is the final relative error in the approximation of the log-determinant. For each value of $\ell$ that we consider, we run each algorithm $100$ times; the plots report the mean relative error, with respect to the target quantity ${\rm trace}\log(A+I)$, and the standard deviation obtained discarding the best and the worst $10\%$ errors over the runs, represented by vertical bars. To make the bars relative to the standard deviation visually distinguishable among the methods, some lines are slightly shifted to the left and to the right, although the budget of matvecs used for different methods is still the same. The plots in Figure 6 show that if we choose a value of $\beta$ that is too small, the method can fail to detect when it is convenient to switch the strategy, and this is due to the fact that the rank-$\beta\ell$-approximation is very different from the rank-$\ell$ approximation. Specifically, in (Example Alg), and (Example Matérn) with $\nu=\frac{1}{2}$, Algorithm 1 with the values $\beta=\frac{1}{8}$, $\beta=\frac{1}{4}$ in (Example RBF) wrongly applies the one-sample strategy. Besides, in (Example RBF), values of $\beta=\frac{1}{8}$, $\beta=\frac{1}{4}$ and $\beta=\frac{1}{2}$ wrongly predicts the mixed strategy. From now on we fix $\beta=\frac{3}{4}$.